What quantity of a 40% acid solution must be mixed with a 10% solution to produce 300 quarts of a 30% solution?
Answer by stanbon(75887)

What quantity of a 40% acid solution must be mixed with a 10% solution to produce 300 quarts of a 30% solution?
----
Equation:
acid + acid = acid
0.40x + 0.10(300-x) = 0.30*300
---
Multiply thru by 100 to get:
40x + 10*300 - 10x = 30*300
30x = 20*300
x = 200 quarts (amt. of 40% solution needed for the mixture)
===============
Cheers,
Stan H.